OpenStudy (anonymous):

what is (11^-4)^-6=11

OpenStudy (anonymous):

does anyone know?

OpenStudy (anonymous):

it is false

Nnesha (nnesha):

the question is \[\huge\rm (11^{-4})^{-6}=11^{??}\] apply the exponent rule \[\huge\rm (a^m)^n=a^{m \times n}\]

OpenStudy (anonymous):

noo, i mean what what would be after 11

Nnesha (nnesha):

here is an exxample \[\huge\rm (x^3)^2=x^{3 \times 2}=x^{6}\]

Nnesha (nnesha):

just multiply the exponents.

Directrix (directrix):

 Use @Nnesha example. Post what you get, okay? @eviequinn111